OpenStudy (anonymous):

Differentiate using chain rule: y = √(2x) * sin*(3x)

OpenStudy (anonymous):

\[y = \sqrt{2x} \times \sin (3x)\]

OpenStudy (baldymcgee6):

\[y = \sqrt{2x}* \sin(3x)\]\[y = (2x)^{1/2} * \sin(3x)\]\[y' = (2x)^{-1/2}*\sin(3x) + (2x)^{1/2}*\cos(3x)*3\]\[y' = \frac{ \sin(3x)+ 3\sqrt{2x} *\cos(3x)}{ \sqrt{2x} }\]

OpenStudy (anonymous):

when you do the 3rd step, in product rule why didn't you make it 1/2(2x)^(-1/2) ??

OpenStudy (baldymcgee6):

y = (2x)^(1/2) y' = 1/2(2x)^-(1/2) * 2 You have to multiply by the derivative of the inside, y ' = (2x)^(-1/2) The 2's cancel

OpenStudy (anonymous):

oh yeah, thanks so much!

OpenStudy (baldymcgee6):

you're welcome